# From Complex-Analytic Models to Sparse Domination: A Dyadic Approach of Hypersingular Operators via Bourgain’s Interpolation Method

### v4

@@ 4. Critical line estimates for the hypersingular Bergman projection
which can be regarded as a singular integral counterpart of the hypersingular maximal operator $M_{t}^{{\mathcal{D}}}$. We first recall that the off-critical line $L^{p}$ theory was studied in [14, Theorem 3]. In particular, they showed that for $K_{2t}$ is bounded from $L^{p}({\mathbb{D}})$ to $L^{q}({\mathbb{D}})$ if and only 1/q1/p>2t21/q-1/p>2t-2 for $1\leq p,q\leq\infty$, which is exactly the off-critical line regime for ${\mathcal{M}}_{t}^{{\mathcal{D}}}$. Therefore, it is natural to consider the behavior of $K_{2t}$ on the critical line
$\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\frac{1}{q}-\frac{1}{p}>2t-2\right\}.$
@@
where in the last estimate, we used the standard integral estimate [37, Theorem 1.12]. Therefore, for any $\alpha>0$, as in Lemma 3.4,
@@
Our next goal is to establish estimates at the other endpoint $\big(\tfrac{1}{3-2t},1\big)$. First, we observe the following unboundedness result.
@@
Let $1<t<3/2$. Then the hypersingular Berezin transform $K_{2t}^{+}$ fail to be bounded from $L^{\frac{1}{3-2t}}({\mathbb{D}})$ to $L^{1,\infty}({\mathbb{D}})$, where we recall that
$K_{2t}^{+}f(z):=\int_{{\mathbb{D}}}\frac{f(w)}{|1-z\overline{w}|^{2t}}\,dA(w).$
@@
Take any $\varepsilon>0$ sufficiently small such that $3-2t-\varepsilon>0$, and let $f(z):=\frac{1}{(1-|z|^{2})^{3-2t-\varepsilon}},\;z\in{\mathbb{D}}$. First, by a direct computation, we have
$\left\|f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}})}\simeq\frac{1}{\varepsilon^{3-2t}}.$
On the other hand, by [37, Theorem 1.12], one has
$K_{2t}^{+}f(z)$ $=\int_{{\mathbb{D}}}\frac{1}{(1-|w|^{2})^{3-2t-\varepsilon}|1-z\overline{w}|^{2t}}dA(w)$
$=\int_{{\mathbb{D}}}\frac{(1-|w|^{2})^{2t+\varepsilon-3}}{|1-z\overline{w}|^{2t}}dA(w)$
$\simeq\frac{1}{(1-|z|^{2})^{1-\varepsilon}}.$
Therefore, for any $\alpha>0$, the above estimate yields
$\alpha\left|\left\{z\in{\mathbb{D}}:\left|K_{2t}^{+}f(z)\right|\gtrsim\alpha\right\}\right|$ $\simeq\alpha\left|\left\{z\in{\mathbb{D}}:\frac{1}{(1-|z|^{2})^{1-\varepsilon}}>\alpha\right\}\right|$
$\simeq\alpha\cdot\alpha^{-\frac{1}{1-\varepsilon}}=\alpha^{\frac{\varepsilon}{1-\varepsilon}}.$
Thus, if $K_{2t}^{+}:L^{\frac{1}{3-2t}}({\mathbb{D}})\to L^{1,\infty}({\mathbb{D}})$ is bounded, then for any $\alpha>0$, one has
$\alpha^{\frac{\varepsilon}{1-\varepsilon}}\lesssim\frac{1}{\varepsilon^{3-2t}},$
which is impossible by letting $\alpha\to\infty$. ∎
Now we turn to some positive results. As discussed in Section 1 (see the comments following (1.4)), the main difficulty is the lack of strong-type bounds along the critical line 1/q1/p=2t21/q-1/p=2t-2. To overcome this difficulty, we combine a “hypersingular” sparse domination with Bourgain’s interpolation trick. (see, e.g., [3, 6]; see also [7] for an abstract extension of this method developed in the setting of fairly general normed vector spaces). The following formulation is the one needed for our purposes.
Lemma 4.3 (Bourgain’s interpolation trick).
$\left\|T_{j}\right\|_{L^{p_{1}}({\mathbb{D}})\to L^{q_{1}}({\mathbb{D}})}\leq M_{1}2^{\beta_{1}j}$
and
$\left\|T_{j}\right\|_{L^{p_{2}}({\mathbb{D}})\to L^{q_{2}}({\mathbb{D}})}\leq M_{2}2^{-\beta_{2}j},$
$\left\|T\right\|_{L^{p,1}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})}\leq CM_{1}^{\theta}M_{2}^{1-\theta},$
where
and $C$ depends only on $\beta_{1}$ and $\beta_{2}$.
We are ready to state the main result in this section.
Theorem 4.4.
For any $1<t<3/2$, $K_{2t}:L^{\frac{1}{3-2t},1}({\mathbb{D}})\to L^{1,\infty}({\mathbb{D}})$ is bounded.
Proof.
Let $f$ be any measurable function on $\mathbb{D}$}. We first observe that $K_{2t}f$ admits a pointwise bound by a finite sum of “hypersingular” sparse operators. This follows from standard dyadic analysis on $\mathbb{D}$} (see, e.g., [15, 18, 31]). For the reader’s convenience, we include a brief sketch of the argument.
First, recall that for any $z,w\in{\mathbb{D}}$, there exists a Carleson tent $Q_{J}$, containing both $z$ and $w$ such that $|Q_{J}|\simeq\left|1-z\overline{w}\right|^{2}$ (see, e.g., [31, Section 3]). Then using the $1/3$-trick, one can find some $I\in{\mathcal{D}}$ or $\widetilde{{\mathcal{D}}}$ (see, Remark 1.4), such that
$J\subset I\quad\textrm{and}\quad|I|\lesssim|J|,$
which further gives
$z,w\in Q_{I}\quad\textrm{and}\quad|Q_{I}|\simeq|1-z\overline{w}|^{2}.$
As a conseqeunce, we have the following pointwise “hypersingular” sparse bound
(4.1) $\left|K_{2t}f(z)\right|=\left|\int_{{\mathbb{D}}}\frac{f(w)}{(1-z\overline{w})^{2t}}dA(w)\right|\lesssim{\mathcal{A}}^{t}_{{\mathcal{S}},{\mathcal{D}}}f(z)+{\mathcal{A}}^{t}_{{\mathcal{S}},\widetilde{{\mathcal{D}}}}f(z),\quad z\in{\mathbb{D}},$
where ${\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}$ is the hypersingular sparse operator associated to a dyadic system $\mathcal{D}$} on $\mathbb{T}$, defined as [^4] We call ${\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}$ a hypersingular sparse operator since, on the critical line $\{(\tfrac{1}{p},\tfrac{1}{q}):\ \tfrac{1}{q}-\tfrac{1}{p}=2t-2\}$, it fails to satisfy strong $(p,q)$ bounds. Indeed, any strong-type estimate for ${\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}$ at a point on this line would imply the corresponding bound for $K_{2t}$, which is known to be false by [14, Theorem 3].
${\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}f(z):=\sum_{I\in{\mathcal{D}}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q_{I}}(z)}{|Q_{I}|^{t}}\int_{Q_{I}}|f(w)|dA(w).$
Therefore, it suffices to prove that $A_{{\mathcal{S}},{\mathcal{D}}}^{t}:L^{\frac{1}{3-2t},1}({\mathbb{D}})\to L^{1,\infty}({\mathbb{D}})$ is bounded. For simplicity, we denote ${\mathcal{A}}_{{\mathcal{S}}}^{t}={\mathcal{A}}^{t}_{{\mathcal{S}},{\mathcal{D}}}$.
Now we shall apply Bourgain’s interpolation trick. To do this, we decompose
${\mathcal{A}}^{t}_{{\mathcal{S}}}=\sum_{j\geq 0}{\mathcal{A}}^{t}_{{\mathcal{S}},j},$
where
${\mathcal{A}}^{t}_{{\mathcal{S}},j}f(z):=\sum_{\begin{subarray}{c}I\in{\mathcal{D}}\\ |I|=2^{-j}\end{subarray}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q_{I}}(z)}{|Q_{I}|^{t}}\int_{Q_{I}}|f(w)|dA(w),\quad z\in{\mathbb{D}}.$
We claim that for any $j\geq 0$,
(4.2) $\left\|{\mathcal{A}}_{{\mathcal{S}},j}^{t}\right\|_{L^{1}({\mathbb{D}})\to L^{1}({\mathbb{D}})}\lesssim 2^{(2t-2)j}$
and
(4.3) $\left\|{\mathcal{A}}_{{\mathcal{S}},j}^{t}\right\|_{L^{\infty}({\mathbb{D}})\to L^{1}({\mathbb{D}})}\lesssim 2^{-(3-2t)j}.$
For the moment, we assume that (4.2) and (4.3) hold. To see the desired restrict weak type bound for ${\mathcal{A}}_{{\mathcal{S}}}^{t}$, we apply Lemma 4.3 with
$\beta_{1}=2t-2,\quad\beta_{2}=3-2t,\quad p_{1}=q_{1}=q_{2}=1,\quad\textrm{and}\quad p_{2}=\infty.$
Therefore,
$\left\|{\mathcal{A}}_{{\mathcal{S}}}^{t}\right\|_{L^{p,1}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})}\lesssim 1,$
for
$\theta=3-2t,\quad\frac{1}{p}=\frac{3-2t}{1}+\frac{2t-2}{p_{2}}=3-2t,\quad\textrm{and}\quad\frac{1}{q}=1.$
This concludes the proof.
Therefore, we only need to show (4.2) and (4.3).
Proof of (4.2). Note that it is more convenient to work with ${\mathcal{A}}_{{\mathcal{S}},j}^{t}$ rather than ${\mathcal{A}}_{{\mathcal{S}}}^{t}$, since the family $\{Q_{I}\}_{I\in{\mathcal{D}},\ |I|=2^{-j}}$ is pairwise disjoint. The argument is straightforward: let $f\in L^{1}({\mathbb{D}})$,
$\left\|{\mathcal{A}}_{{\mathcal{S}},j}^{t}f\right\|_{L^{1}({\mathbb{D}})}$ $=\int_{{\mathbb{D}}}\left(\sum_{\begin{subarray}{c}I\in{\mathcal{D}}\\ |I|=2^{-j}\end{subarray}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q_{I}}(z)}{|Q_{I}|^{t}}\int_{Q_{I}}|f(w)|dA(w)\right)dA(z)$
$=\sum_{\begin{subarray}{c}I\in{\mathcal{D}}\\ |I|=2^{-j}\end{subarray}}\frac{1}{|Q_{I}|^{t-1}}\int_{Q_{I}}|f(w)|dA(w)$
$\simeq 2^{(2t-2)j}\sum_{\begin{subarray}{c}I\in{\mathcal{D}}\\ |I|=2^{-j}\end{subarray}}\int_{Q_{I}}|f(w)|dA(w)$
$=2^{(2t-2)j}\int_{1-2^{-j}\leq|z|<1}|f(w)|dA(w)$
$\leq 2^{(2t-2)j}\left\|f\right\|_{L^{1}({\mathbb{D}})},$
which gives (4.2).
Proof of (4.3). The proof for the second claim is similar. Indeed, for any $f\in L^{\infty}({\mathbb{D}})$, from the above argument, we have
$\left\|{\mathcal{A}}_{{\mathcal{S}},j}^{t}f\right\|_{L^{1}({\mathbb{D}})}$ $\simeq 2^{(2t-2)j}\int_{1-2^{-j}\leq|z|<1}|f(w)|dA(w)$
$\lesssim 2^{(2t-2)j}\cdot\left|\left\{z\in{\mathbb{D}}:1-2^{-j}\leq|z|<1\right\}\right|\cdot\left\|f\right\|_{L^{\infty}({\mathbb{D}})}$
$\simeq 2^{(2t-2)j}\cdot 2^{-j}\cdot\left\|f\right\|_{L^{\infty}({\mathbb{D}})}$
$=2^{-(3-2t)j}\left\|f\right\|_{L^{\infty}({\mathbb{D}})}.$
Hence, (4.3) holds.
The proof of Theorem 4.4 is complete. ∎
Combining Lemma 4.1 and Theorem 4.4 with off-diagonal Marcinkiewicz interpolation theorem, we have the following.
@@
then for every $0<r\leq\infty$, the operator $K_{2t}^{{\mathcal{D}}}$ extends to a bounded map
$K_{2t}^{{\mathcal{D}}}:L^{p,r}({\mathbb{D}})\to L^{q,r}({\mathbb{D}}),$
In particular, $K_{2t}^{{\mathcal{D}}}:L^{p}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})$ is bounded.
Remark 4.6.
Using Bourgain’s interpolation trick, one can give an alternative proof of Corollary 3.7. Indeed, the following pointwise bound is immediate:
${\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)\leq{\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}f(z),\quad z\in{\mathbb{D}}.$
The desired conclusion then follows by repeating the proof of Theorem 4.4 and applying the off-diagonal Marcinkiewicz interpolation theorem.
@@ 5. Hypersingular sparse operators
A natural question arising from the proof of Theorem 4.4 is whether Bourgain’s interpolation trick can be used to study other dyadic operators in harmonic analysis. It turns out that a more general principle underlies this phenomenon.
@@
We begin with the following model case. Let $t>1$, and for simplicity, assume that $\mathcal{S}$ be an $\eta$-sparse family of cubes in $\mathbb{R}^{n}$ such that $Q\subseteq Q_{0}=[0,1]^{n}$ for all $Q\in{\mathcal{S}}$. We define the (pointwise) hypersingular sparse operator associated with $\mathcal{S}$ by
@@
This global containment assumption is natural in the present hypersingular regime. Indeed, as explained in Remark 1.5 (2), when $t>1$ the geometry of the ambient domain becomes relevant. On the other hand, from the viewpoint of sparse domination in harmonic analysis, the assumption is mild: most operators of interest (such as Calderón–Zygmund operators and Hilbert transforms along monomial curves) are local, or can be decomposed into a sum of localized pieces, and this yields the existence of such a global cube $Q_{0}$ (for each localized piece).
@@
Let $\eta\in(0,1)$, $t>1$ and $\mathcal{S}$ be an $\eta$-sparse family in $\mathbb{R}^{n}$ defined as above. For which pairs $1\leq p,q\leq\infty$ does $\mathbb{A}_{\mathcal{S}}^{t}$ extend to a bounded operator
$\mathbb{A}_{\mathcal{S}}^{t}:L^{p,1}(\mathbb{R}^{n})\to L^{q,\infty}(\mathbb{R}^{n})$
(or satisfy a stronger estimate at the pair $(p,q)$, e.g. a weak-type, or strong-type bound, and more generally a Lorentz-space bound)?
@@ 5.1. Graded family
The key observation is already contained in Example 1.9, which shows that one must control how the cube sizes in $\mathcal{S}$} change from one “layer” to the next. This motivates us to introduce the notion of a graded family of dyadic cubes.
@@
We have the following result.
@@
    (Critical line estimate) We consider two different cases.
    1. (a)
      $\mathbb{A}_{{\mathcal{S}}}^{t}$ extends to a bounded operator from $L^{p,\infty}(\mathbb{R}^{n})$ to $L^{q}(\mathbb{R}^{n})$ when $(p,q)$ belongs to
      (5.2) $\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\frac{1}{q}-\frac{1}{p}=\frac{nK_{{\mathcal{S}}}(t-1)}{-\log_{2}(1-\eta)}\quad\textrm{and}\quad q\neq 1\right\}.$
    2. (b)
      $\mathbb{A}_{{\mathcal{S}}}^{t}$ extends to a bounded operator from $L^{p,1}(\mathbb{R}^{n})$ to $L^{q,\infty}(\mathbb{R}^{n})$ when
      (5.3) $p=\frac{-\log_{2}(1-\eta)}{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}\qquad\textrm{and}\qquad q=1.$
@@
Theorem 5.7 strengthens our earlier results on the boundedness behavior of $K_{2t}$ (see, Section 4). Moreover, the proof of Step II in Theorem 5.7 yields an alternative proof of [14, Theorem 3]. Finally, we note that the endpoint weak-type estimate (5.3) fails in general (see, Proposition 4.2).
@@
(5.4) $\sum_{Q\in{\mathcal{S}}^{(j)}}|Q|\lesssim(1-\eta)^{j},$
We divide the proof into several steps.
@@
[…]
By the sparseness and contracting property of $\mathcal{S}$}, $\mathbb{A}_{{\mathcal{S}}}^{t}f(z)$ can only take larger values on $\bigcup_{Q\in{\mathcal{S}}^{(1)}}Q$, whose size is at most $1-\eta$.
@@
$\mathbb{A}_{{\mathcal{S}}}^{t}f(z)\leq 1+\frac{1}{|Q|^{t}}\int_{Q}|f(x)|dx\leq 1+|Q|^{1-t}\leq 1+2^{nK_{{\mathcal{S}}}(t-1)},\quad z\in[0,1]^{n}\backslash\left(\bigcup_{Q\in{\mathcal{S}}^{(2)}}Q\right)$
and $\mathbb{A}_{{\mathcal{S}}}^{t}f(z)$ can only take larger values on $\bigcup_{Q\in{\mathcal{S}}^{(2)}}Q$, whose size is at most $(1-\eta)^{2}$.
@@
where $C_{1},C_{2}>0$ are some absolute constants [^6] Here, we can take $C_{1}=\frac{2^{nK_{{\mathcal{S}}}(t-1)}}{2^{nK_{{\mathcal{S}}}(t-1)}-1}$ and $C_{2}=\frac{1}{2^{nK_{{\mathcal{S}}}(t-1)}-1}$. that only depend on $n,K_{{\mathcal{S}}}$ and $t$, then $z\in[0,1]^{n}\backslash\left(\bigcup_{Q\in{\mathcal{S}}^{(J+1)}}Q\right)$, whose size is at least $1-(1-\eta)^{J+1}$. Therefore, for any $\alpha>0$ sufficiently large, if $\mathbb{A}_{{\mathcal{S}}}^{t}f(z)>C_{1}\alpha-C_{2}$, then
$z\in\bigcup_{Q\in{\mathcal{S}}^{(\widetilde{J}+1)}}Q$
for $\widetilde{J}>\frac{\log_{2}\alpha}{nK_{\mathcal{S}}(t-1)}$, which implies

### v5

@@ 4. Critical line estimates for the hypersingular Bergman projection
which can be regarded as a singular integral counterpart of the hypersingular maximal operator ${\mathcal{M}}_{t}^{{\mathcal{D}}}$. We first recall that the off-critical line $L^{p}$ theory was studied in [14, Theorem 3]. In particular, they showed that for $K_{2t}$ is bounded from $L^{p}({\mathbb{D}})$ to $L^{q}({\mathbb{D}})$ if and only 1/q1/p>2t21/q-1/p>2t-2 for $1\leq p,q\leq\infty$, which is exactly the off-critical line regime for ${\mathcal{M}}_{t}^{{\mathcal{D}}}$. Therefore, it is natural to consider the behavior of $K_{2t}$ on the critical line
$\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\frac{1}{q}-\frac{1}{p}=2t-2\right\}.$
@@
where in the last estimate, we used the standard integral estimate [39, Theorem 1.12]. Therefore, for any $\alpha>0$, as in Lemma 3.4,
@@
Our next goal is to establish weak-type estimates at the other endpoint $\bigl(\tfrac{1}{3-2t},1\bigr)$. We have the following result.
@@
For any $1<t<3/2$, $K_{2t}:L^{\frac{1}{3-2t}}({\mathbb{D}})\to L^{1,\infty}({\mathbb{D}})$ is bounded.
@@
Let $f\in L^{\frac{1}{3-2t}}({\mathbb{D}})$. Write
$K_{2t}f(z)=W(z)B_{2t}f(z),$
where $W(z):=(1-|z|^{2})^{2-2t}$ and $B_{2t}$ is the Forelli–Rudin type operator given by
$B_{2t}f(z):=(1-|z|^{2})^{2t-2}\int_{{\mathbb{D}}}\frac{f(w)}{(1-z\overline{w})^{2t}}dA(w).$
We make the following claims.
  1. (1)
    $B_{2t}:L^{\frac{1}{3-2t}}({\mathbb{D}})\to L^{\frac{1}{3-2t}}({\mathbb{D}})$ is bounded;
  2. (2)
    $W\in L^{\frac{1}{2t-2},\infty}({\mathbb{D}})$.
Assuming (1) and (2), and using Lorentz–Hölder inequality, we have
$\left\|K_{2t}f\right\|_{L^{1,\infty}({\mathbb{D}})}$ $=\left\|W\cdot B_{2t}f\right\|_{L^{1,\infty}({\mathbb{D}})}\lesssim\left\|W\right\|_{L^{\frac{1}{2t-2},\infty}({\mathbb{D}})}\left\|B_{2t}f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}})}$
$\lesssim\left\|W\right\|_{L^{\frac{1}{2t-2},\infty}({\mathbb{D}})}\left\|f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}})}\lesssim\left\|f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}})},$
which gives the desired result.
Therefore, it remains to verify claims (1) and (2). Claim (1) follows from [35, Theorem 3] by choosing $a=2t-2,b=0,c=2t,n=1,\alpha=\beta=0$, and $p=q=\frac{1}{3-2t}$. Claim (2) follows from a direct computation. ∎
Remark 4.3.
The proof of the above proposition illustrates what we will call the Forelli–Rudin method: one reduces the desired weak-type estimate to a weight multiplied by a less singular Forelli–Rudin type operator. While the proof of Proposition 4.2 is carried out in a complex-analytic way, we will see later that the same idea can be adapted to obtain endpoint estimates for certain hypersingular averaging operators (see, Theorem 5.7).
Combining Lemma 4.1 and Proposition 4.2 with off-diagonal Marcinkiewicz interpolation theorem, we have the following.
@@
then for every $0<r<\infty$, the operator $K_{2t}$ extends to a bounded map
$K_{2t}:L^{p,r}({\mathbb{D}})\to L^{q,r}({\mathbb{D}}),$
In particular, $K_{2t}:L^{p}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})$ is bounded.
@@ 5. Hypersingular sparse operators
A natural question arising from the study of ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and $K_{2t}$ is whether there exists a unified harmonic-analytic framework that treats these two complex-analytic models simultaneously, and that extends to a broader class of hypersingular operators. It turns out that a more general principle underlies this phenomenon.
@@
We begin with the following model case. Let $t>1$, and for simplicity, let $\mathcal{S}$ be an $\eta$-sparse family of cubes in $\mathbb{R}^{n}$ such that $Q\subseteq Q_{0}=[0,1]^{n}$ for all $Q\in{\mathcal{S}}$. We define the (pointwise) hypersingular sparse operator associated with $\mathcal{S}$ by
@@
This global containment assumption is natural in the present hypersingular regime. Indeed, as noted in Remark 1.5(2), when $t>1$ the operator $\mathbb{A}_{{\mathcal{S}}}^{t}$ may not even be well-defined on nonzero constant functions whenever it is not localized. On the other hand, from the viewpoint of dyadic harmonic analysis, the assumption is mild: most operators of interest (such as Calderón–Zygmund operators and Hilbert transforms along monomial curves) are local, or can be decomposed into a sum of localized pieces, and this yields the existence of such a global cube $Q_{0}$ (for each localized piece).
@@
Let $\eta\in(0,1)$, $t>1$, and $\mathcal{S}$ be an $\eta$-sparse family in $\mathbb{R}^{n}$ as above. For which pairs $(p,q)$ with $1\leq p,q\leq\infty$ does $\mathbb{A}_{\mathcal{S}}^{t}$ extend to a bounded operator
$\mathbb{A}_{\mathcal{S}}^{t}:L^{p}(\mathbb{R}^{n})\to L^{q}(\mathbb{R}^{n})\qquad\text{or}\qquad\mathbb{A}_{\mathcal{S}}^{t}:L^{p}(\mathbb{R}^{n})\to L^{q,\infty}(\mathbb{R}^{n})?$
Our goal in the remainder of this section is to address this question.
@@ 5.1. Graded family
The key observation is already contained in Example 1.9, which shows that one must control how the sizes of cubes in $\mathcal{S}$} change from one “layer” to the next. This motivates us to introduce the notion of a graded family of dyadic cubes.
@@
We have the following result, whose proof relies on a dyadic version of the Forelli–Rudin method.
@@
    (Critical line estimate) $\mathbb{A}_{{\mathcal{S}}}^{t}$ extends to a bounded operator from $L^{p,\infty}(\mathbb{R}^{n})$ to $L^{q}(\mathbb{R}^{n})$ when $(p,q)$ belongs to
    (5.2) $\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\frac{1}{q}-\frac{1}{p}=\frac{nK_{{\mathcal{S}}}(t-1)}{-\log_{2}(1-\eta)}\right\}.$
@@
Theorem 5.7 strengthens our earlier results on the boundedness behavior of $K_{2t}$ (see, Section 4). Moreover, the proof of Step II in Theorem 5.7 yields an alternative proof of [14, Theorem 3], and the proof of Step III in in Theorem 5.7 offers a different proof of Proposition 4.2.
@@
(5.3) $|\mathfrak{D}_{j}|=\sum_{Q\in{\mathcal{S}}^{(j)}}|Q|\lesssim(1-\eta)^{j},$
where we write $\mathfrak{D}_{j}:=\bigcup_{Q\in{\mathcal{S}}^{(j)}}Q$. We divide the proof into several steps.
@@
$\mathbb{A}_{{\mathcal{S}}}^{t}f(z)\leq\frac{1}{\left|[0,1]^{n}\right|^{t}}\int_{[0,1]^{n}}|f(x)|dx\leq 1,\qquad z\in[0,1]^{n}\backslash\mathfrak{D}_{1}.$
By the sparseness and contracting property of $\mathcal{S}$}, $\mathbb{A}_{{\mathcal{S}}}^{t}f(z)$ can only take larger values on $\mathfrak{D}_{1}$, whose size is at most $1-\eta$.
@@
$\mathbb{A}_{{\mathcal{S}}}^{t}f(z)\leq 1+\frac{1}{|Q|^{t}}\int_{Q}|f(x)|dx\leq 1+|Q|^{1-t}\leq 1+2^{nK_{{\mathcal{S}}}(t-1)},\quad z\in[0,1]^{n}\backslash\mathfrak{D}_{2}$
and $\mathbb{A}_{{\mathcal{S}}}^{t}f(z)$ can only take larger values on $\mathfrak{D}_{2}$, whose size is at most $(1-\eta)^{2}$.
@@
where $C_{1},C_{2}>0$ are some absolute constants [^6] Here, we can take $C_{1}=\frac{2^{nK_{{\mathcal{S}}}(t-1)}}{2^{nK_{{\mathcal{S}}}(t-1)}-1}$ and $C_{2}=\frac{1}{2^{nK_{{\mathcal{S}}}(t-1)}-1}$. that only depend on $n,K_{{\mathcal{S}}}$ and $t$, then $z\in[0,1]^{n}\backslash\mathfrak{D}_{J+1}$, whose size is at least $1-(1-\eta)^{J+1}$.
Therefore, for any $\alpha>0$ sufficiently large, if $\mathbb{A}_{{\mathcal{S}}}^{t}f(z)>C_{1}\alpha-C_{2}$, then $z\in\mathfrak{D}_{\widetilde{J}+1}$, for $\widetilde{J}>\frac{\log_{2}\alpha}{nK_{\mathcal{S}}(t-1)}$, which implies